Find maximum value of y(x) = integral from 0 to x (t+sint).dt in [-1,1]

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student

y(x)=integration(t+sint)dt
=[(t^2)/2 -cos(t)](between 0 to x)
=(x^2)/2 -cos(x)+1
max of function will occur when cox(x) is minimum
=>y(x)max =1/2-cos(1)+1=3/2-cos(1)
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty